40m QIL Cryo_Lab CTN SUS_Lab TCS_Lab OMC_Lab CRIME_Lab FEA ENG_Labs OptContFac Mariner WBEEShop
  40m Log  Not logged in ELOG logo
Message ID: 3558     Entry time: Sat Sep 11 22:42:07 2010
Author: valera 
Type: Update 
Category: PSL 
Subject: PSL update 

- The PMC REFL PD was moved from the temporary location to the one called for by the PSL layout (picture attached). The leakage beams were dumped.

- The FSS reference cavity was aligned using temporary periscope and scanned using NPRO temperature sweep. The amplitude of the sweep (sine wave 0.03 Hz) was set such that the PMC control voltage was going about 100 V p-p with. With rough alignment the visibility was as high as 50% - it will be better when the cavity is locked and better aligned but not better than 80% expected from the mode astigmatism that Tara and I measured on Thursday. The astigmatism appear to come from the FSS AOM as it depends on the AOM drive. We reduced the drive control voltage from 5 V to 4V beyond that the diffraction efficiency went below 50%. The FSS REFL PD was set up for this measurement as shown in the attached picture. There is also a camera in transmission not shown in the picture.

Attachment 1: DSC_2502.JPG  2.652 MB  | Show | Hide all | Show all
Attachment 2: DSC_2505.JPG  1.395 MB  | Hide | Hide all | Show all
DSC_2505.JPG
ELOG V3.1.3-